A stone is thrown from a point A with speed 27 m/s at an angle of 26 degree below the horizontal. The point A is 25 m above the horizontal ground. Find the horizontal range of the stone in meter. Give your answer with one decimal place.


1
Expert's answer
2022-02-01T09:57:33-0500

Solution

Speed in horizontal direction

"u_x=27\times cos26°= 24.27m/s\\u_y=27\times sin26°=11.84m/s"


Height from plane

H=-25m

So using second equation of motion

"-25=11.84t-4.9t^2"

"t^2-2.42t-5.1=0"

SO solutions of this quadratic equation is that

"t=\frac{2.42±\sqrt{5.86+20.4}}{2}\\=\frac{2.42±\sqrt{26.26}}{2}\\=\frac{2.42±5.12}{2}"

Using plus sign

t =3.77sec.

So horizontal range is

R="24.27\times3.77=91.50m"